OpenStudy (anonymous):

(5 points) 3. Olaf has 250 files saved on an external hard drive. Of those files, 12% are photos. (a) Write a percent proportion to find the number of files that are photos. (b) How many photos does Olaf have saved on his external hard drive?

OpenStudy (anonymous):

Plz help will give medal and fan

OpenStudy (anonymous):

Im doing (a) first

OpenStudy (anonymous):

i think the answer is 12% over 250

sammixboo (sammixboo):

Well, a percent proportion is this \(\rm\dfrac{part}{whole}=\dfrac{percent}{100}\)

OpenStudy (anonymous):



sammixboo (sammixboo):

Do you know what our part, whole, and percent is? If there is none for it, then say unknown

OpenStudy (anonymous):

12 is the percent

sammixboo (sammixboo):

Just for know, because we will be solving for the unknown value

OpenStudy (anonymous):

250 is the whole

sammixboo (sammixboo):

Right, so \(\rm\dfrac{part}{whole}=\dfrac{12}{100}\)

sammixboo (sammixboo):

Yes, so \(\rm\dfrac{part}{250}=\dfrac{12}{100}\)

OpenStudy (anonymous):

\[12 / 100=0.12\]

sammixboo (sammixboo):

Do you know how to solve this?

sammixboo (sammixboo):

Good so far :P

sammixboo (sammixboo):

\(\rm\dfrac{part}{250}= 0.12\)

OpenStudy (anonymous):

idk the part how do i get the part

sammixboo (sammixboo):

You just solve it like an equation. Multiply both sides of the equation by 250 \(\rm\dfrac{part}{250}\times250=0.12\times 250\) \(\rm part=0.12\times 250\)

OpenStudy (anonymous):

30

sammixboo (sammixboo):

Correct, so \(\rm part = 30\)